The forecast for Sunday was 10 C but luck was on our side and we had 12 -14 C and lovely sunshine,  so it was decided to carry on and go through the colonies for the first time, with special attention to number 7 as it had a very high ( 50 in a week) varroa count.   As a first measure we will try dusting with icing sugar to encourage bees to groom themselves and clear mites at the same time.
Temperature :  12 -14 C,    Sunny,  little wind.
Hive#1   Inspected by Brian Herbert
Brood box :   14 x 12 ,    over 3 frames of brood at all stages, stores ok, room ok, varroa count 4.  temper 10.
Queen excluder and super added,  some pepper-potting with drone cells needs an eye kept on. may mean the queen is losing her fertility.
Hive#2    Inspected  by Brian Herbert
Brood box:  14 x 12  and half,   1 frame in Deep, 2 frames of brood in half.  stores, room ok.  Queen excluder added , under super put back on top.  varroa count 30,  so dusted with icing sugar.   Temper 10
Hive#3    Inspected by Mel.
Brood box:  Deep National.  Small colony, only half a frame of brood, no stores, needed feeding 50:50 sugar solution added.   varroa count  0,   Temper 10
Hive#4   Inspected by Mel.
Brood box:  Deep National,  4 frames of brood, room ok, no stores, needs feeding. 50:50 sugar solution added.    Qx added under super moved from under brood box. Varroa count 12,   Temper 10.
Hive#5   Inspected by Mel.
Brood box:   Deep National,   3 frames of brood.  room ok,  no stores, 50:50 added.  Qx added under super, moved up from under brood box.  Drone brood and 1 drone seen, 3 mouldy super frames swapped.
Hive#6   Inspected by  Liz.
Brood box:  Deep National.  Brood all stages (BAS)  Q seen -still blue. stores and room ok. varroa count 3.  Temper 8 mouse dropping again on vb.
Hive#7   Inspected by Liz.
Brood box:    All Halves (supers)  Peter Streets Method.    6 -7 frames of B A S.  stores ok, room ok.  Big colony.  Varroa count 30 but last week was 50, dusted frame tops and bees with icing sugar.   also a few chalkbrood.
Hive#8   Inspected by Liz.
Brood box:  National Deep.  Q seen,still blue.  brood ok, stores and room ok.  Under super moved up above Qx and another super added.  varroa count 5.   Temper 9
Hive#9   Warre Hive   Quick check only, light stores so 50:50  feed put on.
Hive#10  Inspected by Richard/Brian
Top Bar Hive,   one extra bar put in main chamber,  small colony but stores and room ok. 50:50 syrup put on to help.
Hive 15 F   Inspected by Liz.
Brood box :   National Deep.  Q seen, still blue.  B A S.  stores and room ok.    varroa count  3,  Temper 8
Hive 15 H   Inspected by Liz.
Poly Nuc.  Making up nicely.   1 drawn frame moved next to brood frames.
 
Richard, Liz and Mel.  Apiary Custodians